# Quantum-Enhanced LLM Efficient Fine Tuning

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) enables efficient fine-tuning of pre-trained language models via low-rank matrix approximation, which is effective in many scenarios. However, its low-rank representation capacity is constrained in complex tasks or high-rank dependency settings, potentially limiting model adaptability. Addressing the expressive bottleneck of classical low-rank approximation in fine-tuning large language models, this paper proposes a parameter-efficient fine-tuning method based on a Quantum Weighted Tensor Hybrid Network (QWTHN), which leverages Quantum Neural Network (QNN). The study investigates quantum-classical hybrid parameter-efficient fine-tuning in low-rank spaces. QWTHN decomposes pre-trained weights into quantum neural network and tensor network representations, utilizing quantum state superposition and other methods to break through classical rank limitations. Experiments show that the proposed quantum fine-tuning technique for large models approaches or even surpasses the parameter efficiency of LoRA. On the CPsyCounD and R1-Distill-SFT datasets, QWTHN, compared to classical LoRA, reduces training loss by up to 15% while using 76% fewer parameters, and achieves an 8.4% performance improvement on the CPsyCounD test set. This research not only realizes lightweight and efficient adaptation of quantum resources to billion-parameter models but also validates the practical path of quantum hardware driven by large model tasks, laying the first engineering-ready technical foundation for future quantum-enhanced AGI systems.
@@
Quantum Computing; LLM; Quantum LLM; Fine Tuning
@@ I Introduction
The rapid advancement of large language models (LLMs) has driven innovations in Parameter-Efficient Fine-Tuning (PEFT) to reduce computational overhead while preserving performance. Classical methods such as Low-Rank Adaptation (LoRA) [1] and DoRA [2] operate under the assumption that weight updates during fine-tuning reside in low-rank subspaces, enabling efficient adaptation through trainable low-rank matrices. Similarly, prefix-tuning optimizes task-specific vectors appended to model inputs, mimicking "virtual tokens" to guide downstream tasks without altering core parameters [3]. While effective, these low-rank approximations inherently limit feature representation adaptability, impair convergence in complex tasks, and exhibit sensitivity to rank selection [4, 1, 5].
@@
Despite these advances, existing quantum-inspired frameworks remain largely theoretical, lacking validation on physical quantum hardware. To bridge this gap, we propose the Quantum Weighted Tensor Hybrid Network (QWTHN), which reparameterizes pre-trained layers into quantum tensor hybrid architectures. By synergizing QNN for capturing complex transformations with tensor-based efficiency, our framework achieves superior parameter-efficient fine-tuning while surpassing conventional LoRA in performance. Our main contributions include:
@@
First Implementation of Quantum Computing Inference on Real Machines: We propose a novel quantum large model fine-tuning algorithm, QWTHN, based on a hybrid quantum-classical neural network. The framework leverages the comprehensive learning capabilities of quantum neural networks and quantum tensor networks to achieve parameter-efficient fine-tuning (PEFT). For the first time, this approach enables inference technology on actual quantum computing hardware.
Significant Reduction in Trainable Parameters: The QWTHN framework reduces the number of trainable parameters by 76% compared to LoRA under equivalent conditions, achieving superior parameter efficiency while maintaining model performance.
Efficient Training: By significantly reducing the number of parameters, the model becomes easier to train. In certain scenarios, the convergence speed can be accelerated by 20% in terms of step size (i.e., the rate of loss value change per unit time), enabling the model to achieve superior fitting performance with fewer training steps. This not only mitigates the risk of overfitting but also holds potential for reducing training costs. Furthermore, the quantum tensor network emphasizes the local correlations within input signals, which helps prevent the training process from being trapped in suboptimal local minima.
@@
Enhancing the Expressive Power of Current Low-Rank Fine-Tuning Networks: We demonstrate that QWTHN not only serves as a viable alternative to classical parameter-efficient fine-tuning methods such as LoRA and DoRA, but also exhibits superior performance across a broader spectrum of fine-tuning tasks. Specifically, on the CPsyCounD and R1-Distill-SFT datasets, QWTHN achieves up to a 15% reduction in training loss while significantly decreasing the number of trainable parameters. Moreover, it improves accuracy metrics by up to 20% on small-scale test sets through optimized parameter adaptation mechanisms.
This paper is structured as follows: Section I introduces the research background of large-scale language models, reviews recent advances in quantum-inspired methodologies for model adaptation, and outlines the significance of our work in addressing existing gaps. Section II critically examines classical fine-tuning techniques for large models and explores quantum-based approaches to integrating quantum principles into LLM adaptation frameworks. Section III details our experimental methodology, including the proposed model architecture, presents empirical results with comparative analyses, and discusses the implications of the findings for efficient model fine-tuning. Finally, Section IV summarizes the research contributions and proposes future directions for advancing quantum-augmented LLM technologies.
@@ II Methods
QNN dynamically generate task-adapted weights through quantum entanglement and superposition, enabling finer-grained feature selection. Here, the MPO extracts foundational features, while the QNN focuses on generating optimized weights, achieving optimal computational resource allocation. Quantum circuits can explore a broader solution space [13], leveraging quantum state superposition to evaluate multiple weight combinations simultaneously, thereby circumventing the local optima limitations inherent in classical low-rank models.
@@ II-A Quantum Tensor Network Based on MPO
To more efficiently represent the low-rank weight matrix W in LoRA, we employ a tensor decomposition method called the Matrix Product Operator (MPO) [12]. By reorganizing the elements of matrix W into a higher-dimensional tensor, we obtain a tensor with 2n indices, as show in Fig 1.
Figure 1: MPO decomposition of the weight matrix W. The local tensors $\mathbf{w}^{(k)}$ are represented by filled circles. The hollow circles denote the input and output indices $i_{l}$ and $j_{l}$, respectively, given the input index $i_{k}$ and the output index $j_{k}$, the slice $\mathbf{w}^{(k)}_{j_{k}i_{k}}$ of the local tensor $\mathbf{w}^{(k)}$ is a matrix.
Let the input space X and output space Y have dimensions $N_{x}$ and $N_{y}$, respectively. Through multilinear algebraic transformations, the original weight matrix $W\in\mathbb{R}^{N_{y}\times N_{x}}$ is mapped to a higher-order tensor:
$\mathcal{W}\in\mathbb{R}^{J_{1}\times\cdots\times J_{n}\times I_{1}\times\cdots\times I_{n}},$
the proposed method operates under the following dimensional constraints:
$\prod_{k=1}^{n}I_{k}=N_{x},\quad\prod_{k=1}^{n}J_{k}=N_{y}.$
Adopting a hierarchical index mapping strategy, we reshape the input vector $x\in\mathcal{X}$ into a multidimensional tensor $(i_{1},i_{2},\ldots,i_{n})$, with the output vector corresponding to $(j_{1},j_{2},\ldots,j_{n})$. We then decompose the LoRA weight matrix W into a product of local tensors ${w}^{(k)}$ via tensor factorization:
$\mathcal{W}_{j_{1}\cdots j_{n},i_{1}\cdots i_{n}}=\mathrm{Tr}\left[\prod_{k=1}^{n}\mathbf{w}^{(k)}_{j_{k}i_{k}}\right],$ (1)
@@
$\mathbf{w}^{(k)}_{j_{k}i_{k}}:D_{k-1}\rightarrow D_{k}.$ (2)
Bond dimension parameter $D=\max\{D_{k}\}$ governs the model’s expressive capacity, with its value positively correlated to the quantum entanglement entropy. By establishing a controllable balancing mechanism between model complexity and expressive power, this approach provides a novel technical pathway for lightweight design of large-scale neural networks. For parameter optimization, the total number of trainable parameters satisfies:
@@
when a uniform bond dimension $D_{k}\equiv D$ is adopted, the total parameter count simplifies to:
@@ II-B QWTHN
The core of quantum neural networks lies in their ability to leverage the superposition and entanglement of quantum states, enabling efficient representation of high-dimensional features in Hilbert space. Assuming input data $x\in\mathbb{R}^{n}$ is encoded as a quantum state $|\psi(x)\rangle$, the system evolves into a new quantum state $|\phi(x)\rangle$ after a series of quantum gate operations. By measuring $|\phi(x)\rangle$, a classical output $y\in\mathbb{R}^{m}$ can be obtained. This process can be described by the following formula:
$y=\mathcal{M}(|\phi(x)\rangle),$ (5)
@@
where $\mathcal{M}(\cdot)$ denotes the measurement operation. Although $y$ is classical data, its generation process involves nonlinear transformations of quantum states, enabling the extraction of features that are difficult to capture with classical methods. While classical neural networks can generate non-harmonic functions, QNN excel at fitting truncated Fourier series [15]. Therefore, the quantum-weighted hybrid network not only learns smooth features but also fills non-harmonic gaps through classical neural networks, significantly reducing training error and improving generalization.
To address the limited expressive power of classical neural networks (including MPO neural networks) in low-rank spaces, we perform element-wise multiplication of the classical output $W_{q}$ from the QNN with the output Y from the classical neural network:
$\tilde{O}=W_{q}\odot Y,$ (6)
@@
where $\odot$ denotes the element-wise multiplication operation. The updated output $\tilde{O}$ retains the low-rank structure while incorporating high-dimensional features extracted by the QNN, thereby enhancing the model’s ability to capture complex nonlinear relationships. Through the element-wise multiplication, the QNN output $W_{q}$ provides additional nonlinear features to $Y$. The elements in $Y$ contain high-order feature interaction information related to the input data [16, 17, 18, 19], which is often neglected in low-rank approximations. Therefore, the expressive power of the updated output $\tilde{O}$ is theoretically improved.
Based on the aforementioned MPO and quantum-weighted hybrid network, we construct the final QWTHN, as shown in Fig 2.
Figure 2: QWTHN. The QWTHN leverages the MPO neural network to reduce the parameter count of LoRA. The input vector of LoRA is reshaped and fed into $\mathrm{MPO_{A}}$, where the last dimension of the output is output mpo $\mathrm{output}_{\mathrm{mpo_{A}}}$=128. This output is then transformed by a classical Multilayer Perceptron ($\mathrm{MLP_{A}}$) into output $\mathrm{output}_{\mathrm{MLP_{A}}}$=4, which corresponds to the number of qubits in the subsequent QNN.The output $\mathrm{output}_{\mathrm{MLP_{A}}}$ is input into the QNN, where it undergoes RY angle encoding and variational quantum state transformations. By exploiting the superposition and entanglement of quantum states, the QNN efficiently represents high-dimensional features in Hilbert space. Finally, the expectation values of PauliZ measurements on each qubit yield an output vector $\mathrm{output}_{\mathrm{QNN}}$ of the same size as $\mathrm{output}_{\mathrm{MLP_{A}}}$. The input to the next layer $\mathrm{MLP_{B}}$ is computed as: $\mathrm{input}_{\mathrm{MLP_{B}}}=\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}+\mathrm{output}_{\mathrm{MLP_{A}}}$, where $\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}$ corresponds to Equation (6). The $\mathrm{MLP_{B}}$ then linearly transforms input $\mathrm{output}_{\mathrm{MLP_{B}}}$ into $\mathrm{output}_{\mathrm{MLP_{B}}}$=128, which is fed into $\mathrm{MPO_{B}}$ to produce the final output of the LoRA layer based on QWTHN. All the above dimensions can be set as hyperparameters.
We employ RY rotation angle encoding to reduce the number of encoding layers, thereby enhancing the training efficiency of the QNN and improving the scalability for real quantum hardware execution. By utilizing parameterized two-qubit gates CRZ (Controlled-Rotation-Z), we can adjust parameters to enhance the exploration of the state space. Our design adopts a block structure, where each block consists of continuous nearest-neighbor interactions and one non-local interaction. This modular approach supports performance enhancement through stacking multiple layers, yet achieves near-saturation of expressive power with fewer layers, thus reducing optimization complexity. This design is particularly well-suited for current quantum computing hardware.
@@ III-A Datasets
The datasets used in this study are publicly available, specifically the CPsyCoun[20] Chinese dataset, the R1-Distill-SFT[21] English dataset and the Chinese-DeepSeek-R1-Distill-data-110k[22].
The CPsyCoun Chinese dataset is based on the open-source version available on the HuggingFace platform, released by the CAS-SIAT-XHai team. It is a psychological counseling dialogue dataset containing real-world Chinese dialogue records from counseling scenarios, covering nine common psychological issues such as depression, anxiety, and interpersonal relationships. The dataset spans seven classic theoretical schools in the field of psychological counseling, representing different theoretical frameworks in counseling practice, and is considered a representative branch of technical methods. The data has been anonymized and annotated with multi-turn dialogue roles between counselors and clients, comprising approximately 16,000 dialogue samples. Some data includes fine-grained labels such as counseling strategies and types of emotional support. This dataset is suitable for research on psychological counseling dialogue generation, sentiment analysis, and intervention strategy modeling, providing a benchmark for exploring the application of natural language processing technologies in the mental health domain.
The R1-Distill-SFT dataset is one of the open-source datasets provided by the ModelScope community. It is a structured question bank aimed at K-12 mathematics education, covering areas such as arithmetic, algebra, geometry, and comparison, along with step-by-step solutions. Each data sample includes the question stem, blanks to be filled, multilingual labels (e.g., English and Chinese cn_k12), standardized answers (annotated using LaTeX mathematical symbols), and verification status. The dataset is suitable for model training (e.g., Supervised Fine-Tuning, SFT) and automated evaluation in educational scenarios. It has been filtered based on quality metrics, covering topics ranging from basic operations to medium complexity, and supports research on personalized learning and reasoning ability assessment. The R1-Distill-SFT dataset focuses on supporting tasks such as Knowledge Distillation and Supervised Fine-Tuning (SFT) for models.
The Chinese-DeepSeek-R1-Distill-data-110k is one of the Chinese datasets provided by the ModelScope community. It is a Chinese instruction fine-tuning dataset generated by distilling DeepSeek-R1, containing approximately 110,000 high-quality multi-turn dialogue samples. This dataset is designed to support the fine-tuning and optimization of the DeepSeek-R1 series of models. It covers a wide range of domains, including general dialogue, knowledge-based question answering, code generation, and mathematical reasoning. The dataset is generated using distillation techniques from large language models and undergoes rigorous cleaning and deduplication processes. The data is structured into input-output pairs with contextual annotations, ensuring high diversity in instructions, strong logical coherence, and broad task coverage. It is particularly suitable for training dialogue models capable of understanding and executing complex instructions.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) enables efficient fine-tuning of pre-trained language models through low-rank matrix approximation, achieving effectiveness in many scenarios. However, its representation capacity is constrained in complex tasks or high-rank dependency settings, potentially limiting model adaptability. To overcome the expressive bottleneck in classical low-rank approximation for fine-tuning large language models (LLMs), we propose Quantum Tensor Hybrid Adaptation (QTHA), a parameter-efficient fine-tuning method that integrates a quantum neural network (QNN) with a tensor network. QTHA explores quantum tensor hybrid fine-tuning within low-rank spaces by decomposing pre-trained weights into quantum neural network and tensor network representations, leveraging quantum state superposition to overcome classical rank limitations. Experiments demonstrate that QTHA achieves performance comparable to or surpassing LoRA in parameter-efficient fine-tuning. Compared to LoRA, QTHA reduces trainable parameters by 76% while reducing training loss by up to 17% and improving test set performance by up to 17% within the same training steps. This research not only enables lightweight adaptation of quantum resources to the billion-parameter models but also validates the feasibility of quantum hardware optimization driven by LLM tasks. It establishes the first engineering-ready foundation for future quantum-enhanced Artificial General Intelligence (AGI) systems.
@@
Quantum Computing; LLM; Quantum LLM; Fine-Tuning
@@ I Introduction
The rapid advancement of large language models (LLMs) has driven innovations in parameter-efficient fine-tuning (PEFT) to reduce computational overhead while preserving performance. Classical methods such as Low-Rank Adaptation (LoRA) [1] and Weighted-Decomposed Low-Rank Adaptation (DoRA) [2] assume that weight updates during fine-tuning lie within low-rank subspaces, enabling efficient adaptation via trainable low-rank matrices. Similarly, prefix-tuning optimizes task-specific vectors appended to model inputs, mimicking “virtual tokens” to guide downstream tasks without altering core parameters [3]. While effective, these low-rank approximations inherently limit feature representation adaptability, impair convergence in complex tasks, and exhibit sensitivity to rank selection [1, 2, 4].
@@
Despite these advances, existing quantum-inspired frameworks remain largely theoretical, lacking validation on physical quantum hardware. To bridge this gap, we propose the Quantum Tensor Hybrid Adaptation (QTHA), which reparameterizes pre-trained layers into quantum tensor hybrid architectures. By synergizing QNN for capturing complex transformations with tensor-based efficiency, our framework achieves superior parameter-efficient fine-tuning while surpassing conventional LoRA in performance. Our main contributions include:
@@
First implementation of quantum computing inference for LLM on quantum hardware: We introduce QTHA, a novel quantum-enhanced fine-tuning algorithm for LLM, based on a hybrid quantum-classical neural network architecture. This framework synergistically combines the expressive power of QNN with the efficiency of tensor networks to achieve PEFT. Notably, QTHA represents the first practical implementation of inference technology on quantum hardware, bridging the gap between theoretical quantum machine learning and deployable solutions.
Significant reduction in trainable parameters: The QTHA method demonstrates remarkable parameter efficiency, reducing trainable parameters by 76% compared to LoRA under identical conditions while preserving model performance. This substantial parameter reduction significantly enhances model trainability. Notably, in specific scenarios, QTHA accelerates convergence by 20% in terms of step size (measured by the rate of loss value change per unit time), enabling the model to achieve optimal fitting performance with fewer training iterations. This dual advantage not only reduces overfitting risks but also shows promising potential for lowering computational costs.
@@
Enhancing the performance of LLM fine-tuning: We demonstrate that QTHA not only serves as a viable alternative to classical parameter-efficient fine-tuning methods such as LoRA and DoRA, but also exhibits superior performance across a broader spectrum of fine-tuning tasks. Specifically, QTHA achieves up to a 17% reduction in training loss while significantly decreasing the number of trainable parameters. Moreover, it improves accuracy metrics by up to 17% on small-scale test sets through optimized parameter adaptation mechanisms.
This study employs a systematic framework to advance quantum-enhanced language model research: Section I establishes the innovative potential of this work in addressing domain-critical challenges by tracing the evolutionary trajectory of LLMs and synthesizing advancements in quantum-inspired adaptation methodologies. Section II deconstructs conventional fine-tuning paradigms while proposing a groundbreaking quantum-integrated strategy that architecturally embeds quantum computational principles into LLM adaptation frameworks. Section III designs an innovative hybrid experimental protocol, demonstrating statistically significant benchmarking results through comprehensive comparative analyses to elucidate the mechanism of resource-efficient quantum-enhanced fine-tuning. The concluding Section IV crystallizes dual-impact theoretical-practical innovations in quantum-enhanced fine-tuning of LLMs, charting transformative pathways for next-generation LLM ecosystems.
@@ II Methods
The QNN leverages the dynamic properties of quantum entanglement and superposition to achieve highly nonlinear feature modeling capabilities. In this framework, the MPO is responsible for efficiently extracting abrupt features, while the QNN focuses on effectively learning periodic features[12, 11]. Through the linear combination of the two, this approach aims to overcome the limitations of classical linear layers in feature learning, thereby achieving optimal allocation of computational resources. Furthermore, quantum circuits can explore a broader solution space[13], circumventing the inherent limitations of local optima of classical low-rank models.
@@ II-A Quantum Tensor Network Based on MPO
To efficiently represent the low-rank weight matrix $\mathbf{W}$ in LoRA, we employ a tensor decomposition method called the Matrix Product Operator (MPO) [11]. By reorganizing the elements of matrix $\mathbf{W}$ into a higher-dimensional tensor, we obtain a tensor with $2n$ indices.
Specifically, let the input space $\mathcal{X}$ and output space $\mathcal{Y}$ have dimensions $N_{x}$ and $N_{y}$, respectively. Through multilinear algebraic transformations, the original weight matrix $\mathbf{W}\in\mathbb{R}^{N_{y}\times N_{x}}$ is mapped to a higher-order tensor:
$\mathcal{W}\in\mathbb{R}^{J_{1}\times\cdots\times J_{n}\times I_{1}\times\cdots\times I_{n}},$ (1)
where
$\prod_{k=1}^{n}I_{k}=N_{x},\quad\prod_{k=1}^{n}J_{k}=N_{y}.$ (2)
Adopting a hierarchical index mapping strategy, we reshape the input vector $x\in\mathcal{X}$ into a multidimensional tensor $(i_{1},i_{2},\ldots,i_{n})$, with the output vector corresponding to $(j_{1},j_{2},\ldots,j_{n})$. We then decompose the weight matrix $\mathbf{W}$ into a product of local tensors ${w}^{(k)}$ via tensor factorization:
$\mathcal{W}_{j_{1}\cdots j_{n},i_{1}\cdots i_{n}}=\mathrm{Tr}\left[\mathbf{w}^{(1)}_{j_{1}i_{1}}\mathbf{w}^{(2)}_{j_{2}i_{2}}\cdots\mathbf{w}^{(n)}_{j_{n}i_{n}}\right],$ (3)
@@
$\mathbf{w}^{(k)}_{j_{k}i_{k}}\in\mathbb{R}^{D_{k-1}\times D_{k}}.$ (4)
Here, bond dimension parameter $D=\max\{D_{k}\}$ governs the model’s expressive capacity, with its value positively correlated to the quantum entanglement entropy. By establishing a controllable balancing mechanism between model complexity and expressive power, this approach provides a novel technical pathway for lightweight design of large-scale neural networks. For parameter optimization, the total number of trainable parameters satisfies:
@@
when a uniform bond dimension $D_{k}\equiv D$ is adopted, and the total parameter count simplifies to:
@@ II-B Quantum Tensor Hybrid Adaptation
The core of quantum neural networks lies in their ability to leverage the superposition and entanglement of quantum states, enabling efficient representation of high-dimensional features in Hilbert space. We designed a highly expressive QNN, in which an input vector $x\in\mathbb{R}^{n}$ is assigned to a quantum state $|\psi(x)\rangle$ via angle embedding using $R_{Y}$ gates:
$|\psi(x)\rangle=\bigotimes_{i=1}^{q}R_{Y}(x_{i})|0\rangle,$ (7)
where $q$ denotes the number of qubits.
The quantum state evolves through a parameterized unitary transformation $U(\theta)$:
$|\phi(x)\rangle=U(\theta)|\psi(x)\rangle=\prod_{l=1}^{L}\left(\bigotimes_{i=1}^{q}R_{Y}(\theta_{l,i})\cdot\text{CRZ}(\theta_{l,ij})\right),$ (8)
where $L$ is the number of layers. CRZ represents controlled-Z rotation gates, which introduces entanglement between qubits.
Each qubit is measured through Pauli-Z observables, yielding expectation values:
$y_{i}=\langle\phi(x)|Z_{i}|\phi(x)\rangle\in[-1,1],$ (9)
with the final output $O_{q}=[y_{1},\ldots,y_{m}]$ being a classical vector.
@@
By leveraging quantum state superposition and entanglement, QNN generate nonlinear features in high-dimensional Hilbert space. Although $y$ is classical data, its generation process involves nonlinear transformations of quantum states, enabling the extraction of features that are difficult to capture with classical methods. While MPO neural networks can generate non-harmonic functions, QNN excels at fitting truncated Fourier series [12, 15], by leveraging quantum state superposition and entanglement, QNN generate nonlinear features in high-dimensional Hilbert space.
To address the limitations of classical neural networks in low-rank spaces, we linearly combine the classical output of the QNN, denoted as $O_{q}$, with the output of the classical neural network, denoted as $O_{c}$, $W_{q}$ and $W_{c}$ are the weights for the quantum neural network output and the classical neural network output, respectively, used to adjust the contributions of $O_{q}$ and $O_{c}$:
$\tilde{O}=W_{q}O_{q}+W_{c}O_{c}.$ (10)
@@
The updated output $\tilde{O}$ retains the low-rank feature learning capability while incorporating the high-dimensional features extracted by the QNN, thereby enhancing the model’s ability to model complex nonlinear relationships. Through this linear operation, the QNN’s output $O_{q}$ provides additional nonlinear features to $\tilde{O}$. The elements in $\tilde{O}$ encompass both harmonic and non-harmonic features [16, 17, 18, 19], extracting harmonic features and complementing the classical neural network. As a result, the expressive power of the updated output $\tilde{O}$ is theoretically enhanced.
Based on the aforementioned MPO and quantum hybrid network, we construct the final QTHA, as shown in Fig. 1. QTHA uses MPO to reduce the number of parameters in LoRA. Specifically, the input vector of LoRA is reshaped and fed into $\mathrm{MPO_{A}}$. The output from $\mathrm{MPO_{A}}$ is then processed by a classical multilayer perceptron ($\mathrm{MLP_{A}}$), which transforms it into a representation corresponding to the number of qubits in the subsequent QNN. Within the QNN, the input is first encoded using RY angles, followed by a variational quantum circuit. These operations leverage quantum superposition and entanglement to efficiently encode high-dimensional features. After performing Pauli-Z measurements on each qubit, a vector of the same length as the output of $\mathrm{MLP_{A}}$ is generated. The input for the next layer, $\mathrm{MLP_{B}}$, is computed as a weighted combination of the QNN output and the output from $\mathrm{MLP_{A}}$. Subsequently, $\mathrm{MLP_{B}}$ applies a linear transformation to this combined input, producing a vector that is then passed to $\mathrm{MPO_{B}}$ to generate the final output. All dimensions within this architecture can be configured as hyperparameters for flexibility and optimization. By using parameterized two-qubit gates CRZ (Controlled-Rotation-Z), we can adjust parameters to enhance the exploration of the state space. Our design adopts a block structure, where each block consists of continuous nearest-neighbor interactions and one non-local interaction. This modular approach supports performance enhancement through stacking multiple layers, yet achieves near-saturation of expressive power with fewer layers, thus reducing optimization complexity.
Figure 1: Schematic diagram of the Quantum Tensor Hybrid Adapation (QTHA). QTHA dynamically adjusts feature weights through parameter tuning and outputs a combination of features from MPO and QNN.
@@ III-A Datasets
The datasets used in this study are publicly available, specifically the CPsyCoun [20] Chinese dataset, the R1-Distill-SFT [21] English dataset and the Chinese-DeepSeek-R1-Distill-data-110k (CH-R1-Math) [22].
CPsyCoun dataset is a Chinese psychological counseling dialogue dataset from HuggingFace (by CAS-SIAT-XHai), featuring anonymized, multi-turn dialogues (16K samples) across 9 common issues (e.g., depression, anxiety) and 7 counseling schools. Includes counselor-client role labels, some with fine-grained annotations (e.g., strategies, emotional support). Ideal for counseling dialogue generation, sentiment analysis, and intervention modeling in NLP for mental health.
R1-Distill-SFT dataset is a K-12 math QA dataset with structured questions (arithmetic, algebra, geometry) and step-by-step solutions. Includes question stems, multilingual labels (e.g., English/Chinese), LaTeX-formatted answers, and verification status. Filtered for quality, it supports SFT, knowledge distillation, and automated evaluation in education.
CH-R1-Math dataset is a Chinese instruction-tuning dataset (110K samples) from ModelScope, distilled from DeepSeek-R1. Covers multi-turn dialogues, QA, code generation, and math reasoning. Rigorously cleaned and deduplicated, it features diverse tasks, logical coherence, and contextual annotations. Optimized for fine-tuning models like DeepSeek-R1.